OpenStudy (anonymous):

Find the product. y5 · y3 is it y^8?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Correct..

OpenStudy (theviper):

y^8 Absolute Correct

OpenStudy (anonymous):

\[\large x^a.x^b = x^{a+b}\]

OpenStudy (theviper):

ly*

mathslover (mathslover):

yes it is correct lemme explain you : \[\large{y^5\times y^3=y^{5+3}=y^8}\] Always remember that : \[\large{a^b\times a^c=a^{b+c}}\]